sakurablxssom
sakurablxssom sakurablxssom
  • 02-06-2020
  • Mathematics
contestada

Please help!

You flip a coin twice.

What is the probability of getting tails and then getting heads?

Write your answer as a percentage.


___%

Please help You flip a coin twice What is the probability of getting tails and then getting heads Write your answer as a percentage class=

Respuesta :

imjustasayin
imjustasayin imjustasayin
  • 03-06-2020

Answer: 25%

Step-by-step explanation:

You have a 1/2 chance of landing a heads or tails both times. So, you multiply 1/2 *1/2 = 1/4. 1/4 = 0.25 = 25%.

Answer Link

Otras preguntas

What three religions claim palestine/israel as a holy land? place an asterisk next to the ones that are universalizing religions?
After t years, the amount of money in a savings account that earns simple interest is p + prt, where p is the starting amount and r is the yearly interest rate.
What does 10/2-3(5•0) equal to?
Suppose Raj receives a raise for his hard work and now earns $6.25 per hour how many hours would he need to work to earn $132?
Read the text carefully, then answer the question located beneath: Time to "Bee" Wary, by Gordon Winters Africanized honey bees were created in Brazil in 1957.
The complex of dna and protein that makes up a eukaryotic chromosome is properly called _____. (concept 12.1)
why was the 30 years war fought?
HELP I GOTTA PASS CLASS
2/3x = -8 Show details
Which of the following minerals are especially important to women’s health? a. Calcium and iron b. Sodium and potassium c. Magnesium and zinc d. Selenium an